# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ Abstract
We discuss the possible topological order/topological quantum field theory of different quantum Hall systems. Given the value of the Hall conductivity, we constrain the global symmetry of the low-energy theory and its anomaly. Specifically, the one-form global symmetry and its anomaly are presented as the organizing principle of these systems. This information is powerful enough to lead to a unique minimal topological order (or a small number of minimal topological orders). Almost all of the known experimentally discovered topological orders are these minimal theories. Since this work is interdisciplinary, we made a special effort to relate to people with different backgrounds by providing translations between different perspectives.
@@ 1 Introduction
Universal aspects of the FQHE phenomenon are well understood. They can be captured by continuum topological quantum field theories (TQFT) enriched by a global $\mathrm{U}(1)$ symmetry (corresponding to conservation of electric charge). A mature microscopic theory of the FQHE in the classic setting of two-dimensional electron gases in a magnetic field also exists. Microscopic theories of FQAH and related lattice systems are under development (see [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] for a sampling of recent papers).
@@
In this context, it is interesting to ask how much of the structure of the universal low-energy theory of an FQHE system can be constrained given only the value of the quantized Hall conductivity $\sigma_{H}$.111 Several authors have explored general constraints on the TQFT given measurable data like $\sigma_{H}$ and $c_{-}$, see e.g., [25, 26, 27, 28, 29]. This is often the first (and perhaps only) piece of experimental information available (for example, in the recent discoveries of the FQAH). What can we say about the low-energy TQFT given the Hall conductivity? Answering this question will be our main focus. We will build on a variety of ideas going back to the very first theoretical papers to organize the possible TQFTs with a given Hall conductance.
@@
Along the way, we will identify the minimal topological order (mTO) [^2] See [30, 31] for recent work on minimal topological order in other contexts. - defined to be the one with the smallest number of distinct anyons - for any given value of the Hall conductivity. We will show that there is either a unique or a very small number of mTO theories at a given $\sigma_{H}$. Specifically for $\sigma_{H}=\frac{p}{q}$ with $p,q$ co-prime integers, and $q$ odd, there is a unique minimal theory with $q$ distinct Abelian anyons. For $q$ even, there are four distinct minimal theories with $3q$ anyons ($q$ of which are non-Abelian and the rest are Abelian) and are variants of the familiar Pfaffian quantum Hall states. Interestingly, almost all the observed FQHE states are described by these mTO theories. Thus, knowledge of $\sigma_{H}$ and the criterion of minimality is often enough to fully determine the low-energy TQFT with no reference to a detailed microscopic theory. This may be useful as more FQHE systems are discovered, especially in lattice settings, with or without a weak magnetic field, where the microscopic physics is potentially rather different from the traditional FQHE systems.
@@
One of the beauties of this topic is that many of the results can be derived independently from different perspectives. The topological order that characterizes FQHE states can be described in multiple equivalent ways. One of them is through a Chern-Simons theory [32]. A different perspective is to just focus on the data (such a braiding, fusion, etc.) needed to characterize the anyon excitations of the topological order. In its general form, this leads to a powerful (but abstract) description of the topological order in terms of a (unitary) Modular Tensor Category (MTC) enriched by the global $\mathrm{U}(1)$ symmetry. A related perspective involves keeping track of the full set of gapped anyons and the action of global symmetries on them, as occurs in any specific microscopic setting. Yet another perspective focuses on the generalized symmetries of the IR TQFT and their ’t Hooft anomalies. Finally, through the ‘bulk-edge’ correspondence that relates Chern-Simons theories in 2+1d dimensions to rational conformal field theories in 1+1d dimensions [32, 33, 34], we may fruitfully view quantum Hall phenomena through the lens of the conformal field theory (see, e.g. [35, 36]).
Throughout the course of our discussion, we will make frequent reference to various bosonic, spin, and spinc TQFTs. To make it clear what we mean when we refer to these, we have included Tab. 1 as a helpful reference.
A particular focus will be on a point of view where the existence of Abelian anyon excitations is related to a (generalized) symmetry of the TQFT with an ’t Hooft anomaly [37]. Indeed, it is well appreciated that a focus on symmetries and their anomalies leads to powerful non-perturbative constraints on the possible low-energy behaviors of physical systems. We note that in the underlying microscopic setting (the UV theory), we only have an ordinary global (also known as zero-form) $\mathrm{U}(1)$ symmetry and possibly some other ordinary symmetries (like translation). In contrast, the IR TQFT, at scales below the gap to anyon excitations, has generalized global symmetries with certain anomalies. These new IR symmetries can be standard emergent symmetries or, as we will discuss below, be a result of symmetry transmutation [38]. The presence of these symmetries in the ground states of the TQFT is associated with the existence of excited anyonic quasiparticle states. Our analysis will make use of the generalized global symmetries and anomalies of the IR theory, which are necessarily present in FQH systems.
@@
Table 1: Types of physical systems and their low-energy TQFTs considered in this paper. The physical systems are characterized by specifying the nature of the fundamental charge carriers, which are taken to have charge one. The nature of composite particles formed by binding an integer number of these fundamental charge carriers is indicated in the second column. The last row describes electronic solids. The distinction between these physical systems is captured in the low-energy TQFT through the possible space-time manifolds in which they can be placed. We limit ourselves to orientable spacetimes. In spin (or Fermionic) TQFTs, the spacetime has to be a spin manifold, and the physical answers can depend on the choice of spin structure. A spinc TQFT is a spin TQFT that permits a particular $\mathrm{U}(1)$ enrichment that corresponds to the structure of the allowed local particles as defined in the third row. It can be placed on non-spin manifolds, provided there is a specific background gauge field.
Our results provide powerful tools for numerical studies of the FQHE, aiding in the identification of the TOs realized in exact diagonalization (ED) and Density Matrix Renormalization Group (DMRG) studies.
@@ 1.1 Outline
In Sec. 4, we relate the flux threading argument (reviewed in Sec. 2.2), the fact that the response theory of the FQH system is not globally well-defined, and the existence of a one-form global symmetry with a particular anomaly (reviewed in Sec. 3.2). The anomalous one-form global symmetry is a reformulation of the statement that the system includes Abelian anyon excitations with particular properties. This reformulation allows us to use the powerful machinery of TQFT.
@@
Sec. 7 describes properties of the one-form global symmetry. We review known material and present various extensions of it, which will be needed later.
@@
Several Appendices address more technical details, some of which are known, and others are new. App. A reviews some known examples of TQFTs and introduces our notation. App. B describes how the one-form global symmetry that leads to $\sigma_{H}$ can be embedded in a larger one-form global symmetry of the system. In App. C, we review and extend the different ways a TQFT can be coupled to a background $\mathrm{U}(1)$ gauge field (the electromagnetic field). App. D reviews how the gauging procedure, which is central in our discussion, affects the quantum dimension of the TQFT. And in App. E, we present a construction of a minimal modular extension of some of the premodular TQFTs that we discuss often.
@@ 2.1 Quantum Hall effect
We review [^3] For a good review of quantum Hall physics, see, e.g., [1] the basic phenomenology of the quantum Hall effects (integer, fractional, spin, etc.). Since every quantum Hall effect has a signature in the Hall conductivity, $\sigma_{H}$, we begin by defining it. In practice, $\sigma_{H}$ can be defined either via electrical transport or thermodynamic quantities. Both of these are routinely used in experiments and give the same answer.
@@
where the derivative is taken while preserving the many-body gap. (2.3) is known as the Streda formula.
@@
[…]
@@
In this paper, we will obtain constraints on the low-energy, i.e., Infra-Red (IR), physics of systems with these two starting points; either $\sigma_{H}=p/q$, or translation symmetry with $\nu=p/q$. If $\sigma_{H}$ is rational, very general arguments, reviewed below, demonstrate that the IR theory must support fractionally charged anyon excitations, have associated topological order, and be described by a Topological Quantum Field Theory (TQFT). We will determine the minimal TQFT consistent with the given $\sigma_{H}$. On the other hand, when translation is assumed and a rational $\nu$ is given, we will analyze the global symmetries and anomalies of this microscopic, i.e., Ultra-Violet (UV), system. We consider the structure of IR TQFTs that preserve the UV symmetries and can match the anomalies. We show that the resulting theories are identical to the ones obtained by the consideration of TQFTs consistent with $\sigma_{H}=p/q$.
@@
Lastly, we note that the discussion here can be easily generalized to systems with multiple independent $\mathrm{U}(1)$ symmetries, such as multi-layer QH systems or 2+1d materials with multiple flavor (such as electron spin and/or “valley”) degrees of freedom. We will use multi-layer systems as an example. In this case, the Hall conductivity is a symmetric matrix $\sigma_{H}^{ab}$, defined as
@@
where $\nu^{(a)}$ is the filling factor of the $a$-th layer.
@@ 2.2 Flux threading argument
[…]
If we set $\Delta\Phi=2\pi$ to be a single flux quantum, then after flux threading, the system returns to the initial configuration, possibly up to a large gauge transformation. Therefore, the final state of the adiabatic flux threading must be an eigenstate of the Hamiltonian. Due to adiabaticity, away from the two edges, the system must remain in the local ground state. We can thus equivalently view the final state as nucleating a quasiparticle and its anti-quasiparticle and separating them to the two boundaries.
@@
The flux-threading argument in this section can be easily generalized to the multi-layer case. The $\mathrm{U}(1)$ symmetry of the $a$-th layer leads to a vison quasiparticle denoted as $v_{a}$. Its charge under the $\mathrm{U}(1)$ symmetry of the $b$-th layer is given by
@@
We can also find the spin of $v_{a}$:

### v3

@@ Abstract
We discuss the possible topological order/topological quantum field theory of different quantum Hall systems. Given the value of the Hall conductivity, we constrain the global symmetry of the low-energy theory and its anomaly. Specifically, the one-form global symmetry and its anomaly are presented as the organizing principle of these systems. This information is powerful enough to lead to a unique minimal topological order (or a small number of minimal topological orders). Almost all of the known experimentally discovered topological orders are these minimal theories. Since this work is interdisciplinary, we made a special effort to relate to researchers with different backgrounds by providing translations between different perspectives.
@@ 1 Introduction
Universal aspects of the FQHE phenomenon are well understood. They can be captured by continuum topological quantum field theories (TQFT) enriched by a global $\mathrm{U}(1)$ symmetry (corresponding to conservation of electric charge). Equivalently, they can be understood as topological orders enriched by $\mathrm{U}(1)$ symmetry as a special case of symmetry enriched topological orders (SETs). A mature microscopic theory of the FQHE in the classic setting of two-dimensional electron gases in a magnetic field also exists. Microscopic theories of FQAH and related lattice systems are under development (see [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] for a sampling of recent papers).
(a) A cartoon illustrating the difference between our approach and the usual classification of SETs. (a) Illustration of the usual approach taken to classify SETs. Here, the topological order is already known, and this information, along with the symmetry, is put into the classification. The classification then provides the possible consistent ways of assigning symmetry charges to the anyons. If the symmetry is $\mathrm{U}(1)$, each symmetry assignment will come with a possibly distinct value of $\sigma_{H}$. In contrast, our approach, illustrated in (b), assumes the $\mathrm{U}(1)$ symmetry and the value of $\sigma_{H}$ and returns a list of consistent topological orders. This is relevant experimentally as in this setting the topological order of the system is not already known.
@@
In this context, it is interesting to ask how much of the structure of the universal low-energy theory of an FQHE system can be constrained given only the value of the quantized Hall conductivity $\sigma_{H}$. [^1] Several authors have explored general constraints on the TQFT given measurable data like $\sigma_{H}$ and $c_{-}$, see e.g., [25, 26, 27, 28, 29]. This is often the first (and perhaps only) piece of experimental information available (for example, in the recent discoveries of the FQAH). What can we say about the low-energy TQFT given the Hall conductivity? We note that this is the opposite direction from the usual classification of SETs. There, we already know the topological order, and we merely want to understand how symmetry labels can be assigned to anyons. In this approach, the fractional Hall conductivity is a consequence of the symmetry assignment. Our contribution, illustrated in Fig. 2(a), is the ability to go in the reverse direction. That is, given the value of the fractional Hall conductivity, can we produce a list of consistent topological orders and give recipes, either experimental or numerical, to distinguish between these possibilities? Answering this question will be our main focus. We will build on a variety of ideas going back to the very first theoretical papers to organize the possible TQFTs with a given Hall conductance.
@@
Along the way, we will identify the minimal topological order (mTO) [^2] See [30, 31] for recent work on minimal topological order in other contexts. — defined to be the one with the smallest number of distinct anyons — for any given value of the Hall conductivity. We will show that there is either a unique or a very small number of mTO theories at a given $\sigma_{H}$. Specifically for $\sigma_{H}=\frac{p}{q}$ with $p,q$ coprime integers, and $q$ odd, there is a unique minimal theory with $q$ distinct Abelian anyons. For $q$ even, there are four distinct minimal theories with $3q$ anyons ($q$ of which are non-Abelian and the rest are Abelian) and are variants of the familiar Pfaffian quantum Hall states. We discuss these mTOs in more detail in the summary of our results in Sec. 1.1. Interestingly, almost all the observed FQHE states are described by these mTO theories. Thus, knowledge of $\sigma_{H}$ and the criterion of minimality is often enough to fully determine the low-energy TQFT with no reference to a detailed microscopic theory. This may be useful as more FQHE systems are discovered, especially in lattice settings, with or without a weak magnetic field, where the microscopic physics is potentially rather different from the traditional FQHE systems.
@@
One of the beauties of this topic is that many of the results can be derived independently from different perspectives. The topological order that characterizes FQHE states can be described in multiple equivalent ways. One of them is through a Chern-Simons theory [32]. A different perspective is to just focus on the data (such a braiding, fusion, etc.) needed to characterize the anyon excitations of the topological order. In its general form, this leads to a powerful (but abstract) description of the topological order in terms of a (unitary) modular tensor category (MTC) enriched by the global $\mathrm{U}(1)$ symmetry [33, 34]. A related perspective involves keeping track of the full set of gapped anyons and the action of global symmetries on them, as occurs in any specific microscopic setting. Yet another perspective focuses on the generalized symmetries of the IR TQFT and their ’t Hooft anomalies. Finally, through the “bulk-edge” correspondence that relates Chern-Simons theories in 2+1d dimensions to rational conformal field theories in 1+1d dimensions [32, 35, 36], we may fruitfully view quantum Hall phenomena through the lens of the conformal field theory (see, e.g. [37, 38]).
A particular focus will be on a point of view where the existence of Abelian anyon excitations is related to a (generalized) symmetry of the TQFT with an ’t Hooft anomaly [39]. Indeed, it is well appreciated that a focus on symmetries and their anomalies leads to powerful nonperturbative constraints on the possible low-energy behaviors of physical systems. We note that in the underlying microscopic setting (the UV theory), we only have an ordinary global (also known as zero-form) $\mathrm{U}(1)$ symmetry and possibly some other ordinary symmetries (like translation). In contrast, the IR TQFT, at scales below the gap to anyon excitations, has generalized global symmetries with certain anomalies. These new IR symmetries can be standard emergent symmetries or, as we will discuss below, be a result of symmetry transmutation [40]. The presence of these symmetries in the ground states of the TQFT is associated with the existence of excited anyonic quasiparticle states. Our analysis will make use of the generalized global symmetries and anomalies of the IR theory, which are necessarily present in FQH systems.
1.1 Summary of main results
The goal of this paper is to characterize the possible topological orders that are consistent with a given Hall conductivity. Throughout the course of our discussion, we will make frequent reference to various bosonic, spin, and spinc TQFTs. To make it clear what we mean when we refer to these, we have included Table 1 as a helpful reference.
@@ 1.1 Summary of main results
Table 1: Types of physical systems and their low-energy TQFTs considered in this paper. The physical systems are characterized by specifying the nature of the fundamental charge carriers, which are taken to have charge one. The nature of composite particles formed by binding an integer number of these fundamental charge carriers is indicated in the second column. The last row describes electronic solids. The distinction between these physical systems is captured in the low-energy TQFT through the possible space-time manifolds in which they can be placed. We limit ourselves to orientable spacetimes. In spin (or Fermionic) TQFTs, the spacetime has to be a spin manifold, and the physical answers can depend on the choice of spin structure. A spinc TQFT is a spin TQFT that permits a particular $\mathrm{U}(1)$ enrichment that corresponds to the structure of the allowed local particles as defined in the third row. It can be placed on nonspin manifolds, provided there is a specific background gauge field.
We first review the crucial observation that a fractional $\sigma_{H}$ implies the existence of a special fractionally charged Abelian anyon, which we call the vison. The vison, denoted by $v$, has charge $Q(v)=p/q=\sigma_{H}$, topological spin $h(v)=p/2q=\sigma_{H}/2$, and it determines the fractional $U(1)$ charges of all other anyons via braiding [i.e., $Q(x)\text{ mod }1=B(v,x)$ for any anyon $x$.] The existence of the vison can be argued in many ways: using the flux threading argument (see Sec. 2.2), by showing that the response theory has a one-form symmetry with a particular anomaly (see Sec. 4), by matching the anomalies of the magnetic translations in the UV theory (see Sec. 5.2), or through formal arguments on symmetry enrichment of MTCs (see Appendix C.3.) We present many different approaches on the vison’s existence to help build physical intuition, and to highlight a variety of complementary perspectives.
As the vison is an Abelian anyon, it generates a (one-form) symmetry group, which we denote ${\cal V}^{s,r}=\{1,v,v^{2},\cdots,v^{s-1}\}$. Here $s=q\ell$ is the order of the group, and $r=p\ell$ specifies the anomaly. Note that the order of the group $s$ must be a multiple of $q$ to ensure that $v^{s}=1$, a transparent boson.
Using this information, we can determine the minimal electronic or spinc topological orders given $\sigma_{H}=p/q$. These minimal orders all have the fewest number of anyons in ${\cal V}^{s,r}$. This means that the order of the vison is $s=q$. As mentioned earlier, if $q$ is odd, then there is a unique minimal theory, ${\cal V}^{q,p}$, which has $q$ anyons that are all powers of the Abelian vison. For $q$ even, there are four minimal theories with $3q$ anyons, all variants of the non-Abelian Pfaffian quantum Hall states. We emphasize that this result was not known for $q$ even and has already been useful in numerical studies of the $3/4$ FQHE [43]. We have included Table 2 as a helpful summary of our results. In addition to the experimentally relevant electronic systems we also considered minimal orders in the settings of: 3D topological insulator surface topological orders, bilayer electronic FQHE, and bosonic FQHE.
Spinc, $\sigma_{H}=p/q$
$q$ odd $\mathcal{V}^{q,p}$ $\mathcal{V}^{q,p}$ $q$ $q$
$q$ even $\mathcal{V}^{2q,2p}$ $\mathsf{Pf}^{f}_{q,p,n}$ $4q$ $4q$ ($3q$) for $n$ even (odd)
Bosons, $\sigma_{H}=p/q$
pqpq even $\mathcal{V}^{q,p}$ $\mathcal{V}^{q,p}$ $q$ $q$
pqpq odd $\mathcal{V}^{2q,2p}=\mathcal{A}^{2,2}\boxtimes\mathcal{A}^{q,p+q}$ $\mathsf{Pf}^{b}_{q,p,n}$ $4q$ $4q$ ($3q$) for $n$ even (odd)
Table 2: A summary of our results. This table gives the list of possible minimal TQFTs by quantum dimension $\mathcal{D}^{2}_{{\cal T}}$ and anyon count $N_{{\cal T}}$ for each FQH setup. See Appendix A for a description of the notation. In all the examples, we find that the minimal TQFTs by anyon count are a subset of the minimal TQFTs by quantum dimension; if there are non-Abelian options, as in the case of the odd $n$ theories above, they are precisely these options.
We are further able to 1) develop a systematic algorithm to compile a list of all the possible topological orders that are consistent with a measured Hall conductivity $\sigma_{H}$, and (perhaps) other topological data, and 2) order the allowed topological orders for a given $\sigma_{H}$. We give more details in Sec. 8.
@@ 1.2 Outline
In Sec. 4, we relate the flux threading argument (reviewed in Sec. 2.2), the fact that the response theory of the FQH system is not globally well defined, and the existence of a one-form global symmetry with a particular anomaly (reviewed in Sec. 3.2). The anomalous one-form global symmetry is a reformulation of the statement that the system includes Abelian anyon excitations with particular properties. This reformulation allows us to use the powerful machinery of TQFT.
@@
Section 7 describes properties of the one-form global symmetry. We review known material and present various extensions of it, which will be needed later.
@@
Several appendices address more technical details, some of which are known, and others are new. Appendix A reviews some known examples of TQFTs and introduces our notation. Appendix B describes how the one-form global symmetry that leads to $\sigma_{H}$ can be embedded in a larger one-form global symmetry of the system. In Appendix C, we review and extend the different ways a TQFT can be coupled to a background $\mathrm{U}(1)$ gauge field (the electromagnetic field). Appendix D reviews how the gauging procedure, which is central in our discussion, affects the quantum dimension of the TQFT. And in Appendix E, we present a construction of a minimal modular extension of some of the premodular TQFTs that we discuss often.
@@ 2.1 Quantum Hall effect
We review [^3] For a good review of quantum Hall physics, see, e.g., [1] the basic phenomenology of the quantum Hall effects (integer, fractional, spin, etc.). Since every quantum Hall effect has a signature in the Hall conductivity $\sigma_{H}$, we begin by defining it. In practice, $\sigma_{H}$ can be defined either via electrical transport or thermodynamic quantities. Both of these are routinely used in experiments and give the same answer.
@@
where the derivative is taken while preserving the many-body gap. Equation (2.3) is known as the Streda formula.
@@
Everything we have said up to this point applies to any system with a gap to excited states carrying current. Henceforth, we specialize to the various quantum Hall systems. In these systems the Hall conductivity $\sigma_{H}$ is quantized to rational numbers in units where the electron charge $e=1$, and $\hbar=1$. [^4] It will be a consequence of our later discussion that any system with a gap to all excitations must have $\sigma_{H}$ a rational number. So we have not lost much generality by specializing to quantum Hall systems. Throughout the text we will write $\sigma_{H}=p/q$, where $p$ and $q$ are coprime integers.
@@
In this paper, we will obtain constraints on the low-energy, i.e., infrared (IR), physics of systems with these two starting points; either $\sigma_{H}=p/q$, or translation symmetry with $\nu=p/q$. If $\sigma_{H}$ is rational, very general arguments, reviewed below, demonstrate that the IR theory must support fractionally charged anyon excitations, have associated topological order, and be described by a topological quantum field theory (TQFT). We will determine the minimal TQFT consistent with the given $\sigma_{H}$. On the other hand, when translation is assumed and a rational $\nu$ is given, we will analyze the global symmetries and anomalies of this microscopic, i.e., ultraviolet (UV), system. We consider the structure of IR TQFTs that preserve the UV symmetries and can match the anomalies. We show that the resulting theories are identical to the ones obtained by the consideration of TQFTs consistent with $\sigma_{H}=p/q$.
@@
Lastly, we note that the discussion here can be easily generalized to systems with multiple independent $\mathrm{U}(1)$ symmetries, such as multilayer QH systems or 2+1d materials with multiple flavor (such as electron spin and/or “valley”) degrees of freedom. We will use multilayer systems as an example. In this case, the Hall conductivity is a symmetric matrix $\sigma_{H}^{ab}$, defined as
@@
where $\nu^{(a)}$ is the filling factor of the $a$th layer.
@@ 2.2 Flux threading argument
Consider the system on an annulus $R_{1}\leq r\leq R_{2},\ 0\leq\varphi<2\pi$ where $r$ and $\varphi$ are the polar coordinates. We assume that $R_{2}\gg R_{1}$. Imagine threading a magnetic flux [^5] Later in the paper, we will refer to this as a holonomy and reserve the use of the term ‘magnetic flux’ to situations where the flux penetrates the space occupied by the charged particles. The only exception is that we will use the term “flux threading argument” when we refer to this argument. We also want to point out that this process of continuously changing the holonomy around a nontrivial cycle in space and following how the spectrum evolves is known as spectral flow and has been used extensively in various branches of mathematics and physics. $\Phi(t)$ through the inner hole. This leads to the vector potential $\vec{A}(t)=\Phi(t)\mathbf{\hat{e}}_{\varphi}/(2\pi r)$ on the annulus, where $\mathbf{\hat{e}}_{\varphi}$ is the unit vector along the transverse direction. We assume that the flux threading is slow and the time evolution of the bulk system is adiabatic. The flux threading induces an electric field $\vec{E}=-\mathbf{\hat{e}}_{\varphi}(d\Phi/dt)/(2\pi r)$, driving a Hall current in the inward radial direction $I_{r}=\sigma_{H}(d\Phi/dt)/2\pi$. If the total flux threaded is $\Delta\Phi$, we find that an amount of charge $\sigma_{H}\Delta\Phi/2\pi$ is accumulated near the inner hole. This setup is depicted in Fig. LABEL:fig:current.
(a) Illustration of the flux-threading argument. (a) The setup of the argument; an annulus with inner radius $R_{1}$ and outer radius $R_{2}$. Flux $\Phi$ is threaded through the hole of the annulus, resulting in a radial electric field, $\vec{E}$. This electric field in turn induces a current flow from the outer edge to the inner edge, $\vec{J}$. When $2\pi$ flux is threaded through the hole, the system can be thought of as having returned to its ground state, albeit having accumulated a charge of $\sigma_{H}$ in the inner hole. (b) Display of the braiding of an anyon $a$ around this flux, which will detect its charge. The flux itself can be thought of as an anyon, the vison $v$, whose charge is $\sigma_{H}$ and whose topological spin is $\sigma_{H}/2$.
If we set $\Delta\Phi=2\pi$ to be a single flux quantum, then after flux threading, the system returns to the initial configuration, possibly up to a large gauge transformation. Therefore, the final state of the adiabatic flux threading must be an eigenstate of the Hamiltonian. Because of adiabaticity, away from the two edges, the system must remain in the local ground state. We can thus equivalently view the final state as nucleating a quasiparticle and its anti-quasiparticle and separating them to the two boundaries.
@@
The flux-threading argument in this section can be easily generalized to the multi-layer case. The $\mathrm{U}(1)$ symmetry of the $a$th layer leads to a vison quasiparticle denoted as $v_{a}$. Its charge under the $\mathrm{U}(1)$ symmetry of the $b$th layer is given by
@@
We can also find the spin of $v_{a}$,